OpenStudy (anonymous):

Which Of the following is the real result of the equation below after completing the square and factoring? x^2 - 4x + 2 = 10 A.) (x-2)^2 = 12 B.) (x-2)^2 = 14 C.) (x+2)^2 = 14 D.) (x+2)^2 = 8 Please HELP!

OpenStudy (shinalcantara):

\[x ^{2} - 4x + 2 =10\] by completing the square: \[x ^{2} - 4x + 4 = 10 -2 + 4\] \[x ^{2} - 4x + 4 = 12\] \[(x - 2)^{2} = 12\]

OpenStudy (anonymous):

Thanks!

OpenStudy (shinalcantara):

yw :)